OpenStudy (anonymous):

How do we know that the equation \[x + \sin(x) + \sqrt{x} = \pi\] has a solution?

myininaya (myininaya):

Let \[f(x)=x +\sin(x)+ \sqrt{x}- \pi\] the domain is [0,inf) the function is continous if we can show that for some number a f(a)=negative number and for some number b f(b)=positive number then there must be a c an element of (a,b) such f(c)=0 so if we can show this then that means f has a solution since it would be it crosses the x-axis

myininaya (myininaya):

f(1)=1+sin(1)+1^(1/2)-pi=-1.12 f(pi)=pi+sin(pi)+pi^(1/2)-pi=pi+pi^(1/2)-pi=pi^(1/2) so there is a number c between 1 and pi such that f(c)=0

OpenStudy (anonymous):

Thank you.